# E-MoFlow: Learning Egomotion and Optical Flow from Event Data via Implicit Regularization

### v1

@@ 4.3 Results on Optical Flow Estimation
We conduct comprehensive quantitative benchmarking for optical flow estimation on the MVSEC dataset, comparing our method against four distinct paradigms: Supervised Learning (SL), Self-Supervised Learning (SSL), Unsupervised Learning (USL), and Model-Based (MB) approaches, as detailed in Tab. 2. The evaluation encompasses two temporal configurations: $dt=1$ indicates evaluating optical flow errors by a duration of a one-frame interval, while $dt=4$ indicates a duration of a four-frame interval. Our method achieves state-of-the-art performance in $dt=1$ settings while attaining the second-highest accuracy in $dt=4$ conditions, surpassed only by the most robust supervised learning baseline ADM-Flow [31], which pretrained on large-scale synthesis dataset. The experimental results demonstrate that by introducing implicit spatial-temporal continuity constraints Eq. (2) and differential geometric constraints Eq. (9), our method can unlock the network’s representation capacity in an unsupervised learning manner, thereby predicting more accurate optical flow.
@@
We performed a qualitative comparison with strong baselines (e.g., [17], [32], [18]) on the MVSEC dataset [53]. As shown in Fig.4, the results demonstrate that our method predicts optical flow closer to the ground truth. Note that we visualize optical flow only at pixels where events are triggered. In addition, ConVGRU-EV-FlowNet [32], EV-FlowNet [18] crop input events to resolution of 256x256, whereas our method and MultiCM [17] do not.
@@ 4.4 Results on 6-DoF EgoMotion Estimation
To evaluate the performance of 6-DoF motion estimation, we compared our method with existing approaches on the MVSEC dataset [53], as it provides ground truth camera motion including angular and linear velocities. As shown in Tab. 3, our method achieves state-of-the-art 6-DoF motion estimation performance in both indoor and outdoor scenes. This is primarily attributed to our method’s use of splines to represent camera motion Eq. (6), which provides a strong continuity prior, combined with our designed differential geometric loss Eq. (9). This loss enables optical flow information to offer geometrically meaningful guidance for motion learning without requiring depth priors. It is important to note that methods like AEmin [42], IncEmin [59], and PEME [43] require depth-enhanced events, leading to stronger performance in the the indoor_flying1 sequence. In addition, the ECN method works well for outdoor_day1 sequence, where motion patterns are relatively simple (with only two dominant motion parameters involved) but struggles with indoor complex motions [48].
@@
indoor_flying1 indoor_flying2 indoor_flying3 outdoor_day1 average $\text{RMS}_{\omega}\downarrow$ $\text{RMS}_{v}\downarrow$ $\text{RMS}_{\omega}\downarrow$ $\text{RMS}_{v}\downarrow$ $\text{RMS}_{\omega}\downarrow$ $\text{RMS}_{v}\downarrow$ $\text{RMS}_{\omega}\downarrow$ $\text{RMS}_{v}\downarrow$ $\text{RMS}_{\omega}\downarrow$ $\text{RMS}_{v}\downarrow$ w/ D AEmin [42] 1.38 0.069 4.00 0.677 IncEmin [59] 1.65 0.085 4.05 1.035 PEME [43] 1.05 0.039 2.72 0.598 w/o D ECN [48] 0.70 MultiCM [17] 7.72 0.24 11.50 0.27 9.53 0.31 6.85 5.90 8.900 1.680 Ours 3.44 0.11 5.31 0.12 4.12 0.15 3.38 0.76 4.062 0.285
Figure 4: Qualitative results of optical flow estimation on the MVSEC dataset. We visualize the optical flow under $dt=4$ configuration. The results in the first, second, and third rows correspond to sequences indoor_flying1, indoor_flying2 and indoor_flying3, respectively. Black points in ground truth (GT) flow maps indicate the absence of LiDAR or depth measurements. The color wheel in the top-left corner encodes the direction and magnitude of the optical flow.
@@ Appendix A Network Architecture
Specifically, our network architecture, inspired by NeRF [61], employs 9 fully-connected layers with 256 dimensional hidden units. The first eight layers utilize ReLU activations to enforce a low Lipschitz constant, ensuring smoother responses to input variations [62], [63]. This design suppresses high-frequency features while favoring learning of low-frequency features, aligning with the prior that optical flow exhibits spatial-temporal smoothness [17, 18, 19]. Notably, no activation function (e.g., ReLU or sigmoid) is applied to the output layer, as optical flow inherently spans both positive and negative values. To further stabilize network training, we introduce residual connections between the second layer to the eighth layer and implement skip connections that concatenate the raw input with the activation outputs of fifth layer. The complete architecture is illustrated in 6.
@@ Appendix C Differential Geometric Loss
Appendix D More Qualitative Results
@@
We further provide additional qualitative results. As shown in 7, our method achieves comprehensive 6-DoF motion estimation on the MVSEC dataset [53]. The angular velocity and linear velocity estimated by our approach closely match the ground-truth motion.
@@
We provide additional qualitative comparisons of optical flow estimation between our method and MultiCM [17], the second-best performing baseline. As shown in 9, our approach predicts optical flow with superior continuity and smoothness, validating the effectiveness of our neural implicit optical flow field representation. The color wheel used to visualize optical flow is shown in 8, where different colors encode the magnitude and direction of the optical flow.

### v2

@@ 4.2 Results on Optical Flow Estimation
We conduct comprehensive quantitative benchmarking for optical flow estimation on the MVSEC dataset [64] and DSEC dataset [68], as detailed in Tab. 1 and Tab. 2. The MVSEC benchmarks optical flow estimation over one-frame ($dt=1$) and four-frame ($dt=4$) intervals, respectively. Our method outperforms all existing unsupervised learning methods on the MVSEC dataset. Notably, under the $dt=1$ setting, our method ranks second among all methods, only behind the traditional optimization-based method MultiCM-V2 [16], and outperforms all supervised learning methods. On the DSEC benchmark, our method performs comparably to existing unsupervised learning approaches. However, there remains a performance gap compared to supervised methods. This is primarily because driving scenes involve more complex motion patterns, abrupt illumination changes, and unstructured scene depth, all of which make it more challenging for unsupervised methods to learn accurate optical flow. The experimental results demonstrate that by introducing implicit spatial-temporal continuity constraints Eq. (2) and differential geometric constraints Eq. (9), our method can unlock the network’s representation capacity in an unsupervised learning manner, thereby predicting more accurate optical flow.
Figure 5: Qualitative results of optical flow estimation on the MVSEC dataset. The results in the first, second, and third rows correspond to sequences indoor_flying1, indoor_flying2 and indoor_flying3, respectively.
@@
We performed a qualitative comparison with strong baselines on the MVSEC dataset and DSEC datasets. As shown in Fig.5, the results on the MVSEC datsets demonstrate that our method predicts optical flow closer to the ground truth. Fig. 6 illustrates that our method also achieved good performance on the DSEC dataset and features smooth optical flow visualizations. Note that we visualize optical flow only at pixels where events are triggered.
@@ 4.3 Results on 6-DoF EgoMotion Estimation
To evaluate the performance of 6-DoF motion estimation, we compared our method with existing approaches on the MVSEC dataset [64], as it provides ground truth camera motion including angular and linear velocities. As shown in Tab. 3, our method achieves state-of-the-art 6-DoF motion estimation performance in both indoor and outdoor scenes. This is primarily attributed to our method’s use of splines to represent camera motion Eq. (6), which provides a strong continuity prior, combined with our designed differential geometric loss Eq. (9). This loss enables optical flow information to offer geometrically meaningful guidance for motion learning without requiring depth priors.
@@
Figure 6: Qualitative results of optical flow estimation on the DSEC dataset. The results in the first, second, and third rows correspond to sequences zurich_city_15a, zurich_city_14c and interlaken_01a, respectively.
4.4 Ablation Study
To validate the efficacy of differential geometric constraints Eq. (9), we conducted ablation studies on the MVSEC dataset [64] under $dt=4$ with three configurations: a) no geometric constraints, b) with differential geometric constraints, and c) direct supervised by ground truth motion. As demonstrated in Tab. 4, the differential geometric constraints yield improvements in optical flow estimation, even achieving performance comparable to those supervised with ground truth motion on indoor_flying2 and indoor_flying3. This indicates that our design enhances the geometric plausibility of the estimated optical flow while effectively avoiding convergence to local minima, as visualized in Fig. 3.
Table 4: Ablation studies on differential geometric constraints.
@@ Appendix A Network Architecture
Specifically, our network architecture, inspired by NeRF [69], employs 9 fully-connected layers with 256 dimensional hidden units. The first eight layers utilize ReLU activations to enforce a low Lipschitz constant, ensuring smoother responses to input variations [70], [71]. This design suppresses high-frequency features while favoring learning of low-frequency features, aligning with the prior that optical flow exhibits spatial-temporal smoothness [17, 18, 19]. Notably, no activation function (e.g., ReLU or sigmoid) is applied to the output layer, as optical flow inherently spans both positive and negative values. To further stabilize network training, we introduce residual connections between the second layer to the eighth layer and implement skip connections that concatenate the raw input with the activation outputs of fifth layer. The complete architecture is illustrated in 7.
@@ Appendix C Differential Geometric Loss
Appendix D More Ablation Studies
Table 5: Ablation studies on early stopping strategy.
[…]
Early Stopping Strategy.
To enhance computational efficiency, we employed the early-stopping strategy from [53, 54, 52] on the MVSEC dataset [64] under $dt=4$. Specifically, we set the patience to 45 and the minimum improvement threshold to $1\times 10^{-3}$, applying the early stopping strategy after 300 iterations. The results in Tab. 5 indicate that the training speed was boosted by 2.2x, while the optical flow estimation accuracy showed only a slight 2.67% drop in the EPE. This demonstrates that the strategy significantly reduces training time at the cost of only a minor loss in accuracy, achieving an excellent trade-off.
Appendix E More Qualitative Results
@@
We further provide additional qualitative results. As shown in 8, our method achieves comprehensive 6-DoF motion estimation on the MVSEC dataset [64]. The angular velocity and linear velocity estimated by our approach closely match the ground-truth motion.
@@
For MVSEC datasets [64], We provide additional qualitative comparisons of optical flow estimation between our method and MultiCM [17], the second-best performing baseline. As shown in 10, our approach predicts optical flow with superior continuity and smoothness, validating the effectiveness of our neural implicit optical flow field representation. The color wheel used to visualize optical flow is shown in 9, where different colors encode the magnitude and direction of the optical flow.
@@
For DSEC datasets [68], We provide additional visualization comparisons of optical flow estimation between our method, state-of-the-art unsupervised learning methods, and supervised learning methods on more sequences. The results in Fig. 11 demonstrate that in some scenarios, our method yields visually superior results with smoother optical flow, while in scenes with complex textures and drastic depth changes, it may produce errors at the edges.
Figure 11: More qualitative results of optical flow estimation on the DSEC dataset. Qualitative results of optical flow estimation on the DSEC dataset. The results in the first, second, third, and fourth rows correspond to sequences interlaken_00b, thun_01a, thun_01a, and zurich_city_14a respectively.
Furthermore, we also provide visualization results of the flow field in XYTX-Y-T 3D space on the MVSEC dataset [64]. The results in Fig. 12 indicate that our method exhibits an emergent capability for point tracking, which is the ability to track the movement of points in pixel space.
Figure 12: Point tracking results of flow field on the MVSEC dataset.